The following contingency table represents the voters’ opinion concerning a new tax



reform bill and their level of income.



Levels of Income



Opinion Low Medium High



For 213 203 182



Against 138 110 154



Apply chi-square (χ2



) test to examine whether the opinion of people on new tax bill is



independent from their income level. [Table value of chi- square (χ2



) at 5% level of



significance and for 2 degrees of freedom is 5.99



1
Expert's answer
2022-01-13T17:26:47-0500

The null hypothesis to be tested is,

"H_0:" The tax opinion is independent from the income level

against the alternative,

"H_1:" The tax opinion is not independent from the income level

The expected counts in the 6 possible combinations are derived from the formula,

"E_{ij}={r_i\times c_j\over n}" where "r_i" is the total for row "i", "c_j" is the total for column "j", "n=1000" is the sample size for "i=1,2" and "j=1,2,3"

The expected counts are,

"E_{11}={r_1\times c_1\over 1000}={598\times351\over1000}=209.898"

"E_{21}={r_2\times c_1\over1000}={402\times351\over1000}=141.102"

"E_{12}={r_1\times c_2\over 1000}={598\times313\over1000}=187.174"

"E_{22}={r_2\times c_2\over1000}={402\times313\over1000}=125.826"

"E_{13}={r_1\times c_3\over1000}={598\times336\over1000}=200.928"

"E_{23}={r_2\times c_3\over1000}={402\times336\over1000}=135.072"

The test statistic is given as,

"\chi^2_c=\sum\sum{(O_{ij}-E_{ij})^2\over E_{ij}}"

Therefore,

"\chi^2_c={(213-209.898)^2\over209.898}+{(182-200.928)^2\over200.928}+{(203-187.174)^2\over187.174}+{(154-135.072)^2\over 135.072}+{(138-141.102)^2\over141.102}+{(110-125.826)^2\over 125.826}=7.87821"


The "\chi^2_{0.05,2}=5.99" as given above and the null hypothesis is rejected if "\chi^2_c\gt\chi^2_{0.05,2}".

Since "\chi^2_{c}=7.87821\gt\chi^2_{0.05,2}=5.99", we reject the null hypothesis and conclude that there is no sufficient evidence to show that the opinion of people on new tax bill is independent from their income level at 5% level of significance.